3x+4+1/2(4x-1)=-3/2
Answer by rfer(16322)

3x+4+1/2(4x-1)=-3/2
3/1x+4/1+4/2x-1/2=-3/2
6/2x+8/2+4/2x-1/2=-3/2
10/2x=-10/2
x=-10/2*2/10
x=-20/20
x=-1